OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 9x + 10 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (accessdenied):

We would check the discriminant. D = b^2 - 4ac The value of the discriminant will determine our solutions To have no real solutions, the discriminant will be negative Two identical rational solutions occur when the discriminant is 0 Two different rational solutions occur when we have a positive discriminant that is a perfect square (examples: 4^2 = 16, 7^2 = 49, etc.) Two irrational solutions occur when we have a positive discriminant that is not a perfect square.

OpenStudy (accessdenied):

you can see how these properties work out when looking at the quadratic formula \[ \frac{-b \pm \sqrt{\color{goldenrod}{b^2 - 4ac}}}{2a}~~<- \color{goldenrod}{discriminant}~in~color\] The discriminant is under that square root if the discriminant is negative, then the square root is not real if the discriminant is 0, then the square root becomes 0 and we end up adding and subtracting 0, which is the same number both ways if the discriminant is postive and perfect square, then the square root will disappear and we'll just have a rational number and, if the discriminant is positive and not perfect, then that square root will stick around, which makes an irrational number.